Circular No. 7630

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001bq IN NGC 5534 E. Cappellaro, Osservatorio Astronomico di Padova; and A. Haru-

tyunyan, Byurakan Astrophysical Observatory, report: “A spectrum of SN 2001bq was obtained on May 19.98 UT with the Asiago 1.8-m telescope + AFOSC (range 350–740 nm, resolution 2.5 nm). This shows a blue continuum and broad P-Cyg Hα and Hβ features, indicating that the supernova is of type II near maximum. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity for NGC 5534 of 2633 km/s yields a supernova expansion velocity of ∼ 11 000 km/s.”

T. Matheson, S. Jha, P. Challis and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2001bq, obtained by P. Berlind on May 20.18 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type II supernova near maximum light. The spectrum consists of a very blue continuum with superposed P-Cyg lines of hydrogen and helium. The above NED recession velocity for NGC 5534 yields the supernova’s expansion velocity (derived from the minimum of the Hβ line) as 9200 km/s.

COMET C/2001 A2 (LINEAR) Z. Sekanina, Jet Propulsion Laboratory, writes: “Astrometric offsets

of component A from B reported between Apr. 30 and May 18 (IAUC 7616, 7625, 7627, MPC 42656, MPEC 2001-K14) indicate that the splitting occurred on Mar. 29.9 ± 1.6 UT (thus coinciding with the major outburst) and that the companion’s relative deceleration is 15.2 ± 0.7 units of 10^{−5} solar attraction. Predicted separations and position angles: May 21, 19′′, 110^{o}; 26, 28′′, 115^{o}; 31, 39′′, 121^{o}; June 5, 54′′, 128^{o}; 10, 74′′, 136^{o}; 15, 97′′, 145^{o}; 20, 122′′, 156^{o}. Assuming that the separation of component C (IAUC 7627) from B coincided with the outburst of May 11, the single available offset suggests that C is subjected to a deceleration of ∼ 40 units of 10^{−5} solar attraction. Preliminary predicted ephemeris: May 21, 3′′, 138^{o}; 26, 8′′, 142^{o}; 31, 16′′, 148^{o}; June 5, 29′′, 154^{o}.”

Total visual magnitude estimates: May 16.96 UT, 5.3 (V. A. Buso, Rosario, Argentina, 0.06-m reflector); 17.90, 5.3 (W. Souza, São Paulo, Brazil, 11 × 80 binoculars); 19.89, 5.2 (Souza, 8 × 30 binoculars); 21.44, 5.4 (A. Pearce, Nedlands, Western Australia, 20 × 80 binoculars); 22.46, 5.2 (Pearce). D. A. J. Seargent, Cowra, New South Wales (25×100 binoculars), writes that on May 14.4 UT there was a narrow ion tail, intense for ∼ 1^{o}, traced for ∼ 4^{o} in p.a. 133^{o}; on May 15.4 a diffuse “cloud” was visible, detached from the tail, at ∼ 1^{o}.5 from the coma.

2001 May 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Brian G. Marsden